solve dx/2x(y+z^2) = dy/y(2y+z^2) = dz/z^3


1
Expert's answer
2021-04-29T18:11:29-0400

The correct form of the given equation is:



"\frac{dx}{2x(y+z^2)}=\frac{dy}{2y(y+z^2)}=\frac{dz}{z^3}"

Taking the first two terms


"\frac{dx}{2x(y+z^2)}=\frac{dy}{2y(y+z^2)}\\
\implies \frac{dx}{x}=\frac{dy}{y}"

Integrating both the sides, we have:


"\log x=\log y + \log c_1\\
x=yc_1\\
c_1 = \frac{x}{y} \qquad \cdots (1)"

Taking Last two terms, we have:


"\frac{dx}{2y(y+z^2)}=\frac{dz}{z^3}\\
\implies z^3dy-2y(y+z^2)dz = 0"

Divide both sides by "-z^3y^2"

"-\frac{1}{y^2}\frac{dy}{dz}+\frac{2}{zy}=-\frac{2}{z^3}\\"

Let

"\frac{1}{y} =t \\ \implies -\frac{1}{y^2}\frac{dy}{dz} =\frac{dt}{dz}"

On comparing it with Linear differential equation, we get:


"P= \frac{2}{z} ,Q= \frac{-2}{z^3}"


Integrating factor, I.F.="e^{\int P.dz}=e^{\int \tfrac{2}{z}dz}=e^{2logz}=z^2"


The solution is:


"t\times I.F.=\int I.F.\times Q dz"

"⇒t(z^2 )=∫z^2×\frac{-2}{z^3}dz"

"\\Rightarrow t\\times z^2=-2logz+c_2\\\\\n\u200b"

"⇒c_2= \frac{z^2}{y}+2logz"

The solution of the given equation is:


"\phi(c_1,c_2)=0 \implies \phi(\dfrac{x}{y},\dfrac{z^2}{y}+2logz)=0"